OpenStudy (anonymous):

Use set builder notation to represent the following set. {... -3, -2, -1, 0} {x | x ∈ I, x ≤ 0} {x | x ∈ N, x ≤ 0} {x | x ∈ R, x ≤ 0}

OpenStudy (anonymous):

What is the set I? The set {...,-2,-1,0} is not a set {x|\[x \in N,x \le 0\]} and {x|\[x \in R\], \[x \le 0\]}

OpenStudy (anonymous):

so it cant be the second and third?

OpenStudy (anonymous):

yup, it is a set {x|\[x \in Z, x \le 0\]}

OpenStudy (anonymous):

I don't know what a set I is, for my educational programs, a set I is known as a set R or C